Hello Expert, Can I get help with this home work. Thank you                                 &nb

Hello Expert,

Can I get help with this home work.

Thank you                                                

   Exercise                                                        

Below are data for four scenarios. Scenario 1 is the base scenario and the other 3 scenarios are modifications to the base scenario.                                               

                                                           

                                        Scenario 1               Scenario 2           Scenario 3           Scenario 4

Sales                                $10,000                $20,000                $12,000             $10,000

COGS                                 8,000                     10,000                   6,000                   8,000

Gross Profit                     $2,000                   $10,000                $6,000                  $2,000

                                                           

Average                           $6,000                   $6,000                 $6,000                  $5,000

Inventory

                                                           

                                                           

Required: Answer the questions below each question and use cell references for calculations.                                                           

1. Compute the gross margin percent for each scenario.                                            

                                                           

                                         Scenario 1              Scenario 2           Scenario 3           Scenario 4

Gross Margin Percent                                                 

                                                           

2. Compute the inventory turnover for each scenario.                                                 

                                                           

                          Scenario 1              Scenario 2           Scenario 3           Scenario 4

Inventory

Turnover                                                          

                                                           

                                                           

3. Compute the gross margin return on inventory investment percent for each scenario.                                             

                                                           

                          Scenario 1              Scenario 2           Scenario 3           Scenario 4

GMROI                                              

                                                           

                                                           

4. For Scenarios 2 through 4, explain what change occurred relative to Scenario 1 to cause the                                                           

gross margin return on inventory (GMROI) to change. You need to specifically discuss each scenario change separately in comparison to scenario 1.                                             

Type your response below:                                          

                                                           

                                                                                         

                                                           

5. Explain to management what type of factors influence the gross margin return on inventory investment.                                                   

Included in your explanation, please included at least one outside reference to support your comments.                                                      

Make sure you are providing detailed explanation, which cannot be completed with a one or two sentence response.                                             

Type your response below:                                                    
Answer #1
Scenario scenario 2 Scenario 3 scenario 4
sales $10,000 $20,000 $12,000 $10,000
COGS 8000 10,000 6000 8,000
gross profit 2000 10000 6000 2000
average inventory 6000 6000 6000 5000
1)gross margin percentage = gross profit/sales*100

=2000/1,0000*100= 20%

=10,000/20,000*100=50% =6000/12000*100=50% =2000/10,000*100=20%
2) inventory turnover= COGS/avg inventory =8000/6000=1.3333333 or 1.34 times =10000/6000=1.666666666=or 1.67 times = 6000/6000=1.00 times =8000/5000=1.60 times
3)GMROI=Gross profit/average inventory

=2000/6000*100=33.33333333%

or 33.34%

=10,000/6000*100=166.666666%

or 166.67%

=6000/6000*100=100% =2000/5000*100=40%4)

4) answer

compare scenario 1 with scenario 2

the GMROI of scenario 1 is 33.34% is very less percentage when compared to the scenario 2 of 166.67%. the reason is the gross profit of the scenario is very low compared to the scenario 2.

compare scenario 1 with scenario 3

scenario 3 has higher gross margin return on inventory investment because the gross profit of the scenario 3 is more than the scenario 1

scenario 1 with scenario 4

scenario 4 has also a higher GMROI of 40% compared to the scenario 1 of 34.34% .

5 answer

the Gross Margin return On Inventory Investment explains the relation between the gross margin and the average inventory .

it is calculated by using the following formula GMROI= gross margin / avg inventory*100

factors influencing on gross margin return on inventory investment is as follows

cost of goods sold ( includes the different manufacturing costs )

sales of the enterprise

opening stock and closing stock

current purchases

outside demand for the product

internal costing procedures etc